OpenStudy (anonymous):

Ryan was 8 when his parents invested $4000 in a certificate of deposit that pays 6%. If Ryan leaves the account alone until it reaches $10,000, how old will he be? (Assume that the interest is not compounded annually.)

OpenStudy (anonymous):

10, 000 = 4,000 * .06 * t => t = 10/ ( 4 * .06) = 41 yrs Thus Ryan will be 8 + 41 = 49 yr old